OpenStudy (anonymous):

if any of these are confusing just ask me to draw them 1) square root of 72 2) square root of 49a^8 3) 2 times the square root of 10 times 3 times the square root of 12 4) square root of 14q times 2 times the square root of 14q 5) 4 divided by the square root of 11

OpenStudy (anonymous):

\[\sqrt{72}\]

OpenStudy (anonymous):

Is your teacher looking for an exact number, or a simplified format?

OpenStudy (anonymous):

simplified

OpenStudy (jacobbenvenutty):

8.48 is the sqr root of 72

OpenStudy (anonymous):

Which identifies all the integer solutions of |x| = 14? A. –14 only B. 0 only C. –14 and 14 D. 14 only

OpenStudy (jacobbenvenutty):

D: 14 only -14 is the complete opposite 0 obviously does not equal 14

OpenStudy (anonymous):

it says for questions 1-4 simplify the radical expression

OpenStudy (anonymous):

Let's start with the first one. Square root of 72 is 8.5. What can we multiply to get this? 6√2 should work

OpenStudy (anonymous):

number one was right mossyfish

OpenStudy (jacobbenvenutty):

if your looking for a radical expression than the answer to question 1. is (8) since the definition of a radical expression is the largest perfect square which will divide evenly into the number under your radical sign.

OpenStudy (anonymous):

For #2, same type of thing. \[\sqrt{49a^8} = \sqrt{7^2(a^4)^2} = 7a^4\]

OpenStudy (anonymous):

For #3 \[ 2√10 \times 3√12 = 12√30 \]

OpenStudy (anonymous):

#4 \[√14q \times 2√4q = 4q√14 \]

OpenStudy (anonymous):

#5 \[4√11 = (4/11)√11\] I hope that this makes sense to you, I apologize for not going in depth with the last few problems

OpenStudy (anonymous):

1)6 square root 2 2)7a^4 3)12 square root of 30 4)28q 5)3.32

OpenStudy (anonymous):

all the answers givin from mossyfish were corect everyone here should become his fan i got 100%

OpenStudy (anonymous):

Thanks!

OpenStudy (jacobbenvenutty):

I will become a fan by your command! lol

OpenStudy (anonymous):

ahah lol